### v1

@@ 5.6 Example: the largest $h^{1,1}(X)$ (polytope with $h^{1,1}=491$)
It is naturally interesting to consider the polytopes with the largest values of the Hodge numbers. It has been known for many years that the largest values of the Hodge numbers $h^{1,1},h^{2,1}$ for toric hypersurface Calabi-Yau threefolds are threefolds with $h^{1,1}(X)=491,h^{2,1}(X)=11$ and $h^{2,1}(X)=491,h^{1,1}=11$, respectively Candelas:1997eh; Aspinwall:1997ye. There is a unique polytope (ID v05-394) with $h^{2,1}(X)=491$, and it has a unique elliptic fibration, namely the standard stacking of fiber $F_{10}$ over Hirzebruch $\mathbb{F}_{12}$; it is easily proven that this is in fact the largest $h^{2,1}(X)$ that can be realized in any elliptic Calabi-Yau threefold, and not just the largest possible in toric hypersurfaces Taylor:2012dr.
@@ 5.6.1 The $F_{10}$ Fiber
We see that we have a “standard stacking” of the base over the fiber vertex $(2,3)$; in fact this is the generic elliptic fibration over the unique toric base with $h^{1,1}(B)=194$, constructed in MorrisonTaylorToric. We therefore have
$T=h^{1,1}(B)-1=193$ (131)
tensor multiplets, in line with the upper bound proposed in Kim:2024eoa.
Because this is the generic fibration over the $h^{1,1}(B)=194$ base, the gauge group of the associated compactification is easy to compute: it is exactly the NHC associated to this base, as identified in MorrisonTaylorToric, which straightforwardly reproduces the gauge algebra in Eq. 127.141414One can equally well identify this gauge algebra by identifying appropriate tops in the polytope, or by passing to Weierstrass form and applying the Kodaira classification; these three approaches all give the same gauge algebra, namely the one in 127.
@@ 5.6.2 The $F_{13}$ Fiber
and thus after implementing the blowup by the coordinate transformation $(a_{210},z)\to(u,uz)$, we find that the 1-1 curve hosts an $\mathfrak{sp}(40)$ gauge algebra, with no additional $(4,6)$ points to resolve. We have therefore shown that the gauge algebra is exactly as in Eq. (128), and moreover the intersection numbers of the curves hosting each summand of the algebra are exactly as described in Kim:2024eoa.
@@ 6 Conclusions
In general, on this second theme, the analysis of this paper illustrates that the identification of fibration structures and the geometric/physical language of F-theory provides a powerful approach to understanding and organizing the otherwise potentially bewildering array of possible Calabi-Yau threefolds. Because almost all known Calabi-Yau threefolds are connected by flops to an elliptic or genus-one fibered phase, it appears that the geometric data (and the associated physical F-theory data) characterizing these fibrations, such as the base variety and the list of divisors hosting gauge algebras in the base, is a useful way of capturing the structure of families of threefolds with a common extended Kähler cone. In particular, since elliptically fibered CY threefolds are all connected in a single moduli space by various (tensor, Higgs, matter) transitions, this provides a way of placing the set of threefolds in the KS database into this broader context. A forthcoming work will show how even the small fraction of non-fibered polytopes can generally be connected into this network through a fairly minimal set of transitions jlrw-forthcoming. Understanding the distribution of and complementarity between different fibration structures for CY threefolds in the same family, i.e. related by flops and flips in the extended Kähler cone, provides another interesting opportunity for further insight into this structure; some aspects of this will be described further in rw-forthcoming. Fibration structure also promises to provide new insights into mirror symmetry Huang:2018vup; ot-forthcoming.
@@
For discrete gauge factors, it is worth noting that toric hypersurface constructions produce genus-one fibrations only with $k$-sections for $k=2,3$, since these are the only genus-one types associated with 2D toric reflexive fibers ($F_{1},F_{2},F_{4}$). An important open question is to identify an upper bound on $k$ for more general Calabi-Yau threefold constructions. It is known that complete intersection fibers can give genus-one fibrations with a 4-section Braun:2014qka. While it has been proven that larger values of $k\geq 5$ cannot be realized through toric constructions or complete intersections, $k=5$ Calabi-Yau threefolds have been constructed using more sophisticated methods Knapp:2021vkm. It has been conjectured Oehlmann:2016wsb that the maximum value of $k$ possible is $k=6$. At the same time, in Raghuram:2018hjn, F-theory models were found that appear to admit a physical Higgsing to $\mathbb{Z}_{21}$, in tension with the conjecture that $k\leq 6$. The class of genus-one fibered Calabi-Yau threefolds with $k>4$ represents perhaps the least-understood corner of the space of elliptic and genus-one fibered Calabi-Yau threefolds, and presents important challenges for further research.
We have presented here a range of specific examples from the data provided from the full analysis of fibration structures in the KS database. We expect that there is much more that can be learned from this data. A natural next step would be to apply some similar analysis to fibration structures for complete intersections in toric varieties, which would use as fibers intersections in higher-dimensional reflexive subpolytopes, such as the intersection fibers analyzed in Braun:2014qka.
@@ Acknowledgements.
We would like to thank Lara Anderson, Volker Braun, James Gray, Shing Yan Li, Damien Mayorga Pena, and Paul Oehlmann for helpful discussions. This work was supported by the DOE under contract #DE-SC00012567. WT would like to thank the Santa Fe Institute, where some of this work was carried out. RN is a Pappalardo Postdoctoral Fellow in the MIT Department of Physics. FA was supported by the Margaret MacVicar UROP Memorial Fund for this project.
@@ Appendix A Details of Weierstrass Models
In this appendix we review the Weierstrass forms associated with the 16 reflexive 2D polytope toric fibers, originally analyzed in Klevers:2014bqa; see also an2001jacobians. In principle, the procedure is to plug the defining polynomials for each of the 16 reflexive polytopes into Nagell’s algorithm to find their generic Weierstrass models. However, the defining polynomials of the 16 fibers are closely related, and so in practice this exercise would be highly redundant. For instance, from Figure 1 it is straightforward to see that $F_{6}$ can be obtained from $F_{1}$ by adding a vertex, namely the point (1,1)(-1,1). Adding a point to a polytope removes points from its dual, and hence the defining polynomial of the Calabi-Yau hypersurface defined by $F_{6}$ can be obtained by removing monomials from the polynomial of $F_{1}$, i.e. by setting their coefficients to be zero. Actually, $F_{6}$ can be obtained in a second way, by adding the vertex $(0,1)$ to $F_{4}$; this provides an alternative but equivalent way of deriving the Weierstrass form of the defining polynomial of $F_{6}$.
@@
Each of the sixteen reflexive polytopes can be realized in at least one way by adding points to one of $F_{1}$, $F_{2}$, $F_{4}$, and $F_{10}$, and therefore specializations of Eqs. 139, A, 143, and 145 suffice to specify the defining polynomials of all 16 fiber types. The appropriate specializations have been worked out in e.g. Klevers:2014bqa, and are summarized for in Table 2. From the specializations, one can readily work out the generic gauge groups of Klevers:2014bqa by plugging the forms of $f,g$ into the formula for the discriminant $\Delta$ and factoring it, as described in the main text.
@@ Appendix B Statistics of Polytope Automorphisms
In the course of our work, we enumerated the automorphisms of every polytope in the KS list, using the built-in capabilities of CYTools Demirtas:2022hqf; to the best of our knowledge, this is the first time that automorphisms have been exhaustively and systematically studied. While it is not the primary purpose of our work, we exhibit some statistics of those automorphisms in this Appendix.
Of the 473,800,776 polytopes, only 1.3% (6,095,530 total) have at least one nontrivial automorphism. The polytope with the largest automorphism group is the 24-cell, studied in e.g. Braun:2011hd; Braun:2015jdy. This polytope, which in our convention is labeled as v24-506398, is a self-dual polytope with Hodge numbers $h^{1,1}(X)=h^{2,1}(X)=20$ and has an automorphism group of order 1,152. The symmetry group is the Weyl group of the $F_{4}$ root system, and can be given explicitly as
@@
Group GAP ID GAP4 Order #(Polytopes)
@@
Table 12: The 35 nontrivial groups that appear as the automorphism group of a 4D reflexive polytope. For each group, we list its order and the number of polytopes for which it is the automorphism group. “Gap ID” refers to the entry of the group in the online library of small groups, accessed through the GAP system GAP4.

### v2

@@ 5.6 Example: the largest $h^{1,1}(X)$ (polytope with $h^{1,1}=491$)
It is naturally interesting to consider the polytopes with the largest values of the Hodge numbers. It has been known for many years that the largest values of the Hodge numbers $h^{1,1},h^{2,1}$ for toric hypersurface Calabi-Yau threefolds are threefolds with $h^{1,1}(X)=491,h^{2,1}(X)=11$ and $h^{2,1}(X)=491,h^{1,1}=11$, respectively Candelas et al. (1997); Aspinwall and Morrison (1997). There is a unique polytope (ID v05-394) with $h^{2,1}(X)=491$, and it has a unique elliptic fibration, namely the standard stacking of fiber $F_{10}$ over Hirzebruch $\mathbb{F}_{12}$; it is easily proven that this is in fact the largest $h^{2,1}(X)$ that can be realized in any elliptic Calabi-Yau threefold, and not just the largest possible in toric hypersurfaces Taylor (2012).
@@ 5.6.1 The $F_{10}$ Fiber
We see that we have a “standard stacking” of the base over the fiber vertex $(2,3)$; in fact this is the generic elliptic fibration over the unique toric base with $h^{1,1}_{*}(B)=194$, constructed in Morrison and Taylor (2012c). We therefore have
$T=h^{1,1}_{*}(B)-1=193$ (132)
tensor multiplets, in line with the upper bound proposed in Kim et al. (2024).
Because this is the generic fibration over the $h^{1,1}_{*}(B)=194$ base, the gauge group of the associated compactification is easy to compute: it is exactly the NHC associated to this base, as identified in Morrison and Taylor (2012c), which straightforwardly reproduces the gauge algebra in Eq. 128.161616One can equally well identify this gauge algebra by identifying appropriate tops in the polytope, or by passing to Weierstrass form and applying the Kodaira classification; these three approaches all give the same gauge algebra, namely the one in 128.
@@ 5.6.2 The $F_{13}$ Fiber
and thus after implementing the blowup by the coordinate transformation $(a_{210},z)\to(u,uz)$, we find that the 1-1 curve hosts an $\mathfrak{sp}(40)$ gauge algebra, with no additional $(4,6)$ points to resolve. We have therefore shown that the gauge algebra is exactly as in Eq. (129), and moreover the intersection numbers of the curves hosting each summand of the algebra are exactly as described in Kim et al. (2024).
@@ 6 Conclusions
In general, on this second theme, the analysis of this paper illustrates that the identification of fibration structures and the geometric/physical language of F-theory provides a powerful approach to understanding and organizing the otherwise potentially bewildering array of possible Calabi-Yau threefolds. Because almost all known Calabi-Yau threefolds are connected by flops to an elliptic or genus-one fibered phase, it appears that the geometric data (and the associated physical F-theory data) characterizing these fibrations, such as the base variety and the list of divisors hosting gauge algebras in the base, is a useful way of capturing the structure of families of threefolds with a common extended Kähler cone. In particular, since elliptically fibered CY threefolds are all connected in a single moduli space by various (tensor, Higgs, matter) transitions, this provides a way of placing the set of threefolds in the KS database into this broader context. A forthcoming work will show how even the small fraction of non-fibered polytopes can generally be connected into this network through a fairly minimal set of transitions Anderson et al. . Understanding the distribution of and complementarity between different fibration structures for CY threefolds in the same family, i.e. related by flops and flips in the extended Kähler cone, provides another interesting opportunity for further insight into this structure; some aspects of this will be described further in Nally and Taylor . Fibration structure also promises to provide new insights into mirror symmetry Huang and Taylor (2019b); Oehlmann and Taylor .
@@
For discrete gauge factors, it is worth noting that toric hypersurface constructions produce genus-one fibrations only with $k$-sections for $k=2,3$, since these are the only genus-one types associated with 2D toric reflexive fibers ($F_{1},F_{2},F_{4}$). An important open question is to identify an upper bound on $k$ for more general Calabi-Yau threefold constructions. It is known that complete intersection fibers can give genus-one fibrations with a 4-section Braun et al. (2015b). While it has been proven that larger values of $k\geq 5$ cannot be realized through toric constructions or complete intersections, $k=5$ Calabi-Yau threefolds have been constructed using more sophisticated methods Knapp et al. (2021). It has been conjectured Oehlmann et al. (2016) that the maximum value of $k$ possible is $k=6$. At the same time, in Raghuram and Taylor (2018), F-theory models were found that appear to admit a physical Higgsing to $\mathbb{Z}_{21}$, in tension with the conjecture that $k\leq 6$. The class of genus-one fibered Calabi-Yau threefolds with $k>4$ represents perhaps the least-understood corner of the space of elliptic and genus-one fibered Calabi-Yau threefolds, and presents important challenges for further research.
We have presented here a range of specific examples from the data provided from the full analysis of fibration structures in the KS database. We expect that there is much more that can be learned from this data. A natural next step would be to apply some similar analysis to fibration structures for complete intersections in toric varieties, which would use as fibers intersections in higher-dimensional reflexive subpolytopes, such as the intersection fibers analyzed in Braun et al. (2015b).
@@ Acknowledgements.
We would like to thank Lara Anderson, Volker Braun, James Gray, Shing Yan Li, Damien Mayorga Pena, Paul Oehlmann, and Thorsten Schimannek for helpful discussions. This work was supported by the DOE under contract #DE-SC00012567. WT would like to thank the Santa Fe Institute, where some of this work was carried out. RN is a Pappalardo Postdoctoral Fellow in the MIT Department of Physics. FA was supported by the Margaret MacVicar UROP Memorial Fund for this project.
@@ Appendix A Details of Weierstrass Models
In this appendix we review the Weierstrass forms associated with the 16 reflexive 2D polytope toric fibers, originally analyzed in Klevers et al. (2015); see also An et al. (2001). In principle, the procedure is to plug the defining polynomials for each of the 16 reflexive polytopes into Nagell’s algorithm to find their generic Weierstrass models. However, the defining polynomials of the 16 fibers are closely related, and so in practice this exercise would be highly redundant. For instance, from Figure 1 it is straightforward to see that $F_{6}$ can be obtained from $F_{1}$ by adding a vertex, namely the point (1,1)(-1,1). Adding a point to a polytope removes points from its dual, and hence the defining polynomial of the Calabi-Yau hypersurface defined by $F_{6}$ can be obtained by removing monomials from the polynomial of $F_{1}$, i.e. by setting their coefficients to be zero. Actually, $F_{6}$ can be obtained in a second way, by adding the vertex $(0,1)$ to $F_{4}$; this provides an alternative but equivalent way of deriving the Weierstrass form of the defining polynomial of $F_{6}$.
@@
Each of the sixteen reflexive polytopes can be realized in at least one way by adding points to one of $F_{1}$, $F_{2}$, $F_{4}$, and $F_{10}$, and therefore specializations of Eqs. 140, A, 144, and 146 suffice to specify the defining polynomials of all 16 fiber types. The appropriate specializations have been worked out in e.g. Klevers et al. (2015), and are summarized for in Table 2. From the specializations, one can readily work out the generic gauge groups of Klevers et al. (2015) by plugging the forms of $f,g$ into the formula for the discriminant $\Delta$ and factoring it, as described in the main text.
@@ Appendix B Statistics of Polytope Automorphisms
In the course of our work, we enumerated the automorphisms of every polytope in the KS list, using the built-in capabilities of CYTools Demirtas et al. (2022); to the best of our knowledge, this is the first time that automorphisms have been exhaustively and systematically studied. While it is not the primary purpose of our work, we exhibit some statistics of those automorphisms in this Appendix.
Of the 473,800,776 polytopes, only 1.3% (6,095,530 total) have at least one nontrivial automorphism. The polytope with the largest automorphism group is the 24-cell, studied in e.g. Braun (2012); Braun et al. (2015a). This polytope, which in our convention is labeled as v24-506398, is a self-dual polytope with Hodge numbers $h^{1,1}(X)=h^{2,1}(X)=20$ and has an automorphism group of order 1,152. The symmetry group is the Weyl group of the $F_{4}$ root system, and can be given explicitly as
@@
Group GAP ID 41 Order #(Polytopes)
@@
Table 12: The 35 nontrivial groups that appear as the automorphism group of a 4D reflexive polytope. For each group, we list its order and the number of polytopes for which it is the automorphism group. “Gap ID” refers to the entry of the group in the online library of small groups, accessed through the GAP system 41.